2. A researcher wishes to test the claim of a particular cereal manufacturer that the mean weight of cereal in the boxes is less than 300g. A sample of 50 boxes yields a sample mean weight of 296g. Assume that the population standard deviation is 5g.


a) Can we conclude that the claim is true? Test at α = 0.05.


b) Obtain a 95% confidence interval for μ.

1
Expert's answer
2021-05-03T04:55:23-0400

a) The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq300"

"H_1:\mu<300"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value is "z_c=-1.6449."

The rejection region for this left-tailed test is "R=\{z:z<-1.6449\}."  


The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{296-300}{5/\sqrt{50}}=-5.6569"

Since it is observed that "z=-5.6569<-1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 300, at the 0.05 significance level.


Using the P-value approach:

The p-value for left-tailed test, "z=-5.6569" is "p<0.00001," and since "p<0.00001<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 300, at the 0.05 significance level.


b) The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}},\bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(2.96-1.96\times\dfrac{5}{\sqrt{50}},296+1.96\times\dfrac{5}{\sqrt{50}})"

"=(294.614, 297.386)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "294.614<\mu<297.386," which indicates that we are 95% confident that the true population mean "\mu"  is contained by the interval "(294.614, 297.386)."